### v1

@@ 3 Forming membrane’s potential
We considered that the neuron has a stable base state. On the one side, this resting state must be dynamically stabilized for little perturbations (and to provide a mechanism when the cell grows, divides, or ages). On the other side, it must be able to restore the state after rough perturbations, causing short-time transients (when restoring the membrane’s potential after issuing a spike). The two states need different physical mechanisms.
3.1 Neuron as system controller
From control theory, it is known that the goal of the system is to govern the application of system inputs to drive the system to a desired state while minimizing any delay, overshoot, or steady-state error and ensuring control stability. The neuron implements a controller that monitors the controlled process variable (membrane voltage) and compares it with the reference or setpoint (the resting potential). The difference between the process variable’s actual and desired values called the error signal, is the actual offset potential. It is applied as feedback to generate a control action to bring the controlled process variable to the same value as the set point.
Biology uses such a simple controller. The gradients are used to adjust the process variable by their positive and negative contributions, and the different speeds of the thermodynamical and electrical interactions minimize the delay (i.e., provide the maximum speed of operations, vital for survival). The steady-state error is minimized by setting the process variable to the reference point using long-term stable parameters (geometry and overall concentration). The low-intensity current through the always-open resting ion channels provides dynamic stability in the steady state. The permanently zero error signal indicates a lack of operations. In contrast, the permanently nonzero error signal is a sign of abnormal operation and is likely a symptom of a neurological disease.
The system aims to provide an intense output that informs the downstream neurons. This overshoot is implemented by a large number of voltage-gated ion channels (not identical with the non-gated ion channels used to maintain the resting potential) distributed in the neuron’s wall. The overshoot current flows through a similarly large number of persistently open ion channels concentrated in AIS (it has about two orders of magnitude higher density of channels than that in the wall). The intense slow current produces a condenser-like behavior (capacitive current); the phenomena called polarization and hyperpolarization (not polarization; instead, a movement of completely separated charge in the surface layer of the electrolyte) of the membrane provide the necessary positive and negative error signals for the controller in the transient state by moving the actual potential value of the membrane above and below the resting potential value.
From a control theoretical point of view, the condenser geometry sets the ’always the same’ value (the set-point) and the currents serve to keep or restore that value. Forming membrane’s potential is an excellent example of the principle that nature uses "one process to build and the other to correct" [6].
3.1.1 Biology builds
After introducing a finite-width membrane into an electrolyte, a potential difference between the two membrane surfaces is created; see Eq.(19). When adjusting the membrane’s potential, we must consider its ground and excited states separately, given that the perturbations in these two cases are vastly different. Nature employs various mechanisms to maintain the ground state and recover it after generating a spike. Although nature’s tools are remarkably similar (channels and pumps; here, we discuss only the ’downhill’ channels), the significant difference in current intensity needs a different discussion.
3.1.2 Electricity and thermodynamics correct
The potential’s and the electric field’s magnitudes depend on the concentration and the geometry (the finite width of the membrane). The ions’ chemical nature comes into play only if the polarizability can differ for different molecules.
Given that the same number of charged ions must be present on the two sides of the membrane, their resultant surface density $\sigma_{A}$ must be the same. Notice that the effect is purely electrostatic, resulting in an asymmetric ion distribution; no permeability is needed. If the membrane is (slightly) permeable, ions will move across the membrane until equilibrium is reached. The resulting potential difference is directly proportional to the concentration difference.
The classical condenser has an electric field, as the blue dashed diagram line shows in Fig. 4. No field is outside or inside the condenser plates (see Fig. 3). There is a sudden jump in the field on the conductor’s surface (the armatures) and a constant field between the plates. (The charged layer can be ideally thin if electrons are sitting on the surface but is of finite thickness on the opposite side where ions are on the surface. This latter effect is not discussed in classical electricity.)
The biological condenser behaves differently; see the red diagram line. The electrolyte (instead of an insulator) outside the condenser drastically changes the structure of the field. The electric field on the surface and between the plates changes by a factor of 3 to 4. Outside the plates, the polarization creates a field that changes logarithmically. An ion layer with finite thickness is built near the membrane’s surface. It is uniformly charged, so the field inside it is linear. At the internal surface of the condenser, it takes the value of the field calculated for the gap; on the other side of the layer, it takes the value of the logarithmic field at that position. The dashed line represents the gap field, the continuous line represents the charged layer field, and the dotted line represents the electrolyte field. We can observe the "correspondence principle": the particle-continuous fields join seamlessly; furthermore, if the polarizability of the electrolyte decreases toward zero, the field value approaches the classical value.
The thickness of the charged layer has a significant impact on the resulting electric field (i.e., the neuron’s ’setpoint’). If biological fragments are formed and settle down on the membrane’s intracellular side, due to the increased $\Delta z$, the resting potential (and the threshold potential) may change, leading to neurological diseases.
@@ 3.2 States
The two electric fields show essentially different dependencies on the neuron’s parameters. Recall that although the parameters we used are biologically plausible, they are only estimations. The qualitative conclusions, however, remain valid. As Figure 1 shows, in the function of the membrane’s thickness, the electric field across the membrane due to electric charges (it does not depend on the thickness of the membrane), for different physically plausible assumptions; furthermore, the thermal electric field.
@@
The balanced state is set to where the thermodynamic and electric "electric field" diagram lines cross each other; around $5\ nm$; the value we assumed in evaluating our equations. On the other hand, we can check the electric field’s dependence on the summed-up concentration of the ions in the segment (see Fig. 5). Here, the thermodynamic electric field is constant but different at different width parameters. The classic condenser does not provide a reasonably balanced state (a matching line). The dielectric diagram line seems to provide a realistic estimation when using biologically reasonable parameters. One can describe the neuron as a balanced system at the intersection of the sloping and the horizontal lines. In the resting state, only minor deviations from this state exist in both the electric field and the chemical concentration.
@@
Figure 5: The electric and the thermal "electric fields" at different concentrations, see Eq.(18).The thermal electric field varies with the thickness parameter and depends on the concentration; see Eq.(4). Balanced states are where the sloping lines cross the horizontal ones. The dot marks where HH provided measured data.
@@
In the transient state, if we assume that a charge transfer causes a significant potential change around $100\ mV$ in the layer of width $100\ nm$, it means a $0.1*10^{6}\bigl{[}\frac{V}{m}\bigr{]}$ jump in the electric field that corresponds to a $100\bigl{[}mM\bigr{]}$ jump in the concentration (for only a short time and only in that mentioned layer not in the bulk; see also Fig. 7). During issuing an AP, the system finds its way back to the crossing point (the path is not simple: both concentrations on both sides of the membrane and the electric field in the mentioned layer change; the process is relaxation, as it is seen in the shape of an AP).
@@
Even we can imagine how the electric process happens in this controlled system (for visibility, the positions of the marks are not proportional to the mentioned numbers). The system is in a balanced state at a crossing point. The two circles connected by bent arrows represent that when an excitation begins at the point with lower concentration and higher potential (the cytoplasm side and negative resting potential), ions rush into the surface layer, so the ion concentration increases, that increases the electric field; which is seen that the membrane potential suddenly increases. The increased potential starts a current, but the current is slow, and the intensity of the current through the channels in the membrane’s wall is higher than the current of the drain (AIS) so the system issues an AP while returning to its starting point. Given that it receives a large amount of $Na^{+}$ ions, its primary effort directs to get rid off them (the electric gradient cannot distinguish the electric charges by their chemical nature), so the output current through the AIS fundamentally comprises $Na^{+}$ ions. The system will move along the line, with the restriction that it is a movement in the phase space (the change in concentration is not proportional with the time), and the different points on the membrane’s surface may have the potential at different times (the spatiotemporal behavior). Due to the capacitive current, the local potential may temporarily drop below the set-point (as observed in physiology, the phenomenon is known as hyperpolarization). The time course of the processes are described by the laws of motion of biology [7] (the time derivatives of the Nernst-Planck equation), and the simplified model works as described in section 4 of [30] and its computational details in [31].
@@ 3.2.1 Resting state
In the resting state, the goal is to provide stability by using slow and less intense currents. As [1] discusses, there are ’holes’ (non-controlled ion channels) in the membrane where the counterforce is missing. The difference in the electric and thermodynamic forces could accelerate ions until the Stokes-Einstein speed (see Eq.(27)) is reached. Given that the deviation from the set-point is slight, the intensity and speed of the current are low. The slow ion transport delivers both charge and mass, so the gradients permanently direct the process control variable toward the reference point. Given that the ’resting channels’ are scattered on the surface of the membrane and it takes time for the gradients delivered by slow ions to reach the entrance of an ion channel (see also section 4), some slight variations in the actual value of the membrane’s potential necessarily exist. Given that the AIS comprises non-gated ion channels with much higher density than those in the membrane, the variations and the control ("leaking") are marginal, and for mainly through the ion channels in the membrane’s walls, as observed in experimental physiology.
The electric voltage contribution is unidirectional, the two-component thermodynamic contributions, depending on the concentrations, can be positive or negative. If choosing the zero potential at the middle of the electric potential, we can interpret that the concentration combinations maintain their ionic concentrations (and so also thermodynamic voltage contributions) independently on both sides starting from a zero potential level. They generate potentials autonomously in the segments proximal to the opposite surfaces of the membrane. The electric charge on the membrane creates another voltage independently. However, that voltage must fit between the two thermodynamic voltages. Any deviation in the potentials starts a current directing the concentrations and voltages toward a balanced state. The process is complex: the sum of surface concentration of ions must be the same on both sides for the condenser and the ratios on the intracellular and extracellular sides must be adjusted to satisfy the conditions
@@
In a balanced state, the left sides are zero: at the contact point between the electrolyte and the membrane there are no potential differences, while in perturbed state they are non-zero and so they represent the driving forces for restoring the balance. (Recall that the ions are slow and that the local electric and concentration gradients control the local potentials.) From the derivation, follows that
$U_{electric}$ $+\displaystyle=-2*(U_{internal}^{K^{+}}$$+U_{internal}^{Na^{+}})$ (30)
@@
Figure 6: The mechanism for controlling neuronal operation is the same in all cases shown in Table 1. It appears that the thicker membrane produces a higher membrane electric potential and requires a higher thermal potential to compensate for it. However, it can solve the task using a lower salt concentration. Given that the electric membrane potential accelerates the ions across the ion channels in the membrane, and the operation is faster. The electric field has a similar value, as shown in Fig. 1. Furthermore, the total concentration decreases as the membrane thickness and potential increase.
@@
The membrane’s electric potential defines the resting potential, and the "electric" and "thermodynamic" potentials work together to control neuronal operation. The anatomically defined "electric" membrane potential plus the (compared to the internal amount) vast amount of ions in the extracellular segment and so the stable concentration provides a solid reference point. In the resting state, the membrane potential is balanced the presence of always-open "resting ion channels". When ions flow into the membrane, the "resting ion channels" can keep the balance: the internal $K^{+}$ and $Na^{+}$ concentrations are adjusted as the $Na^{+}$ flows in.
@@
Figure 6 also shows that in its native state, the neuron is electrically balanced: in the internal and external segments, the resulting thermodynamical and electrical forces are equal. Any foreign invasion into the neuron’s life (chemical or electrical, including clamping) changes the left-side reference value to an externally set value. The neuron must adjust its $Na^{+}$ and $K^{+}$ concentrations to reach a new balance while the invasion persists. That is, any invasion ("testing in the physical laboratory") changes the concentrations and voltages inside the cell. In addition, the internal biological processes can trigger internal processes that contribute (from the measurement’s point of view) foreign currents or processes, see also section 3.2.2. The feedback the clamping methods apply introduces a compensating current. Even a simple conductance measurement device is active: it applies a test voltage that generates a test current, which sensitively affects the ionic composition of the neuron (recall that the concentrations may differ by up to six orders os magnitude). Although modern electronic devices attempt to conceal this effect, they subtly influence biological operation.
@@
From our result, it follows that for more chemical elements, a per-element coupled set of equations exists plus the sum concentration agreed electric fields that are valid simultaneously and define the process variable (aka membrane potential) of the neural regulatory circuit. In a more symmetrical form
@@
00 0 $=\frac{U_{electric}^{Total}(c,\Delta z,d)}{2}+\sum_{k}U_{thermal}^{C_{k}}(d)+{invasion}\quad for\ all\ k;\ inside\ and\ outside$ (33)
@@
That is, we have an unidirectional electric potential difference and oppositely directed thermal voltages. In the balanced state, the two potentials must be equal. When changing say one concentration, it changes the other concentration on the same side, and the surface charge density of the condenser. The changed electric voltage re-adjusts the thermodynamic voltage, and as a consequence, the concentrations, in the other segment. There is no simple way to express, say, $[Na^{+}]_{i}$ in the function of $[K^{+}]_{i}$ or vice versa, and similarly for the other ions. Everything moves. The size matters: the amount of ions in the external segment is many orders of magnitude higher, and the gradients are reciprocally proportional to the amount of ions.
As the figure suggests, an external invasion, typically, an electric voltage on the intracellular side, changes the balanced state and due to the parameters are linked, changes all concentrations. When moving the system out of its balanced state, in any way, a driving force appears that moves the system towards finding a new balanced state. However, recall that the ions are slow, so the changes are not instant.
Table 1 shows our results for three different famous published cases. It looks like the calibration of the concentration of negative and positive ions needs scaling and the measurement of $Ca^{2+}$ comprises issues.
@@
$K^{+}$ 400 20 $-75.5=58log\bigl{(}\frac{20}{400}\bigr{)}$ 41.9 41.4 -62.0
$Na^{+}$ 50 440 $+54.8=58log\bigl{(}\frac{440}{50}\bigr{)}$
$Cl^{-}$ 40 560 $-66.5=58log\bigl{(}\frac{560}{40}\bigr{)}$ 41.9 42.0 63.0
$Ca^{2+}$ $0.27^{*}$ 10 $+45.5=29log\bigl{(}\frac{10}{0.27}\bigr{)}$
@@
$K^{+}$ 124 2.25 $-101=58log\bigl{(}\frac{2.25}{124}\bigr{)}$ 88.9 83.6 -125.4
$Na^{+}$ 10.4 109 $+59.2=58log\bigl{(}\frac{109}{10.4}\bigr{)}$
$Cl^{-}$ 1.5 77.5 $-99.4=-58log\bigl{(}\frac{77.5}{1.5}\bigr{)}$ 88.9 82.7 124.1
$Ca^{2+}$ $0.021^{*}$ 2.1 $+58.0=29log\bigl{(}\frac{2.5}{0.25}\bigr{)}$
@@
$K^{+}$ 140 5 $-89.7=62log\bigl{(}\frac{5}{140}\bigr{)}$ 57.7 57.3 -85.8
$Na^{+}$ 15 145 $+61.1=62log\bigl{(}\frac{145}{15}\bigr{)}$
$Cl^{-}$ 4 110 $-89.2=-62log\bigl{(}\frac{110}{4}\bigr{)}$ 57.7 56.9 85.9
$Ca^{2+}$ $0.04^{*}$ 5 $+60.8=31log\bigl{(}\frac{5}{0.04}\bigr{)}$
@@
$Na^{+}$ 12 145 $+66.6=62log\bigl{(}\frac{145}{12}\bigr{)}$ 60.7 66.6

### v2

@@ 3 Forming membrane’s potential
We model the neuron as an electric circle that comprises an electrolyte condenser implemented by its membranes, surrounded by two electrolyte segments. The membrane comprises low-transmission-capacity ion channels in its wall and high-transmission-capacity channels toward its axon; furthermore, it uses slow ions instead of electrons. The condenser produces a potential (that depends on the membrane’s thickness and the concentrations in the two segments, as discussed above). Furthermore, the two electrolyte segments produce two voltages, which are combined from Nernst voltages. According to Kirchhoff’s Voltage Law, in a balanced state, the sum of those voltages must be zero. However, one must take into account the effects of the "slow current" (emulated by "fast current" with a plausible time course). We omit the biological regenerative processes (such as ATP producing ions through hydrolysis, ion pumps and channels functioning, and so on) and discuss only the most important physical primary processes. In light of our findings, we must reconsider several key concepts in neurophysiology.
@@ 3.1 States
The two electric fields show significantly different dependencies on the neuron’s parameters. Recall that although the parameters we used are biologically plausible, they are only estimations. The qualitative conclusions, however, remain valid. Figure 1 shows, in the function of the membrane’s thickness, the electric field across the membrane due to electric charges (it does not depend on the thickness of the membrane), for different physically plausible assumptions; furthermore, the thermal electric field.
@@
The balanced state is set where the thermodynamic and electrical "electrical field" diagram lines cross each other; around 5nm5\ nm; the value we assumed in evaluating our equations. On the other hand, we can check the electric field’s dependence on the summed-up concentration of the ions in the segment (see Fig. 8) (dehydration, for example, changes the setpoint through changing concentrations). Here, the thermodynamic electric field is constant but different at different width parameters. The field of the classic condenser does not provide a reasonably balanced state (a matching line). The dielectric diagram line seems to provide a realistic estimation when using biologically reasonable parameters. One can describe the neuron as a balanced system at the intersection of the sloping and the horizontal lines. In the resting state, only minor deviations (fluctuations) from this state exist in both the electric field and the chemical concentration.
@@
Figure 8: The electric and the thermal "electric fields" at different concentrations, see Eq.(18). The thermal electric field varies with the thickness parameter and depends on the concentration; see Eq.(4). Balanced states are where the sloping lines cross the horizontal ones. The dot marks where $H$ provided measured data.
@@
In the transient state, if we assume that a charge transfer causes a significant potential change around 100mV100\ mV in the layer of width 10nm10\ nm, it means a $10^{7}\bigl[\frac{V}{m}\bigr]$ jump in the electric field that corresponds to a $100\bigl[mM\bigr]$ jump in the concentration (for only a short time and only in that mentioned layer; but not in the bulk; see also Fig. 10). One can infer that the neuron can minimize its offset voltage (and so: its energy) by either decreasing its membrane’s width, the sum concentration of the electrolyte, the electric field, or a process with a temporal course that changes their combination, thereby implementing a true multi-physics phenomenon. During issuing an AP, the system finds its way back to the crossing point (the path is not simple: concentrations on both sides of the membrane and the electric field in the mentioned layers change, all having different speeds; electric, thermodynamical, mechanical and so on, changes happen; the process is a relaxation (a waving), as it is seen in the shape of an AP).
@@
We can imagine how the electric process happens in this controlled system (for visibility, the positions of the marks are not proportional to the mentioned numbers). The system is in a balanced state at a crossing point. The two circles connected by bent arrows represent that when an excitation begins at the point with lower concentration and higher potential (the cytoplasm side and negative resting potential), ions rush into the surface layer, so the ion concentration increases, which increases the electric field; this is seen that the membrane potential suddenly increases. (We just note here, that, due to ions’ repulsion, the pressure also increases, proportionally with the potential.) The increased potential starts a current, but the current is slow, and the intensity of the current through the channels in the membrane’s wall is much lower than the current of the drain (AIS) so the system issues an AP while returning to its starting point. Given that it receives a large amount of $Na^{+}$ ions, its primary effort is directed to get rid of them (the electric gradient cannot distinguish the electric charges by their chemical nature), so the output current through the AIS fundamentally comprises $Na^{+}$ ions. The system will move along the line, with the restriction that it is a movement in the phase space (the change in concentration is not proportional to time), and the different points on the membrane’s surface may have the potential at different times (the spatiotemporal behavior). Due to the capacitive current, the local potential may temporarily drop below the set-point (as observed in physiology, this phenomenon is known as hyperpolarization). The time course of the processes is described by the laws of motion of biology [8] (the time derivatives of the Nernst-Planck equation). The simplified model operates as described in section 4 of [63] and its computational details are provided in [64].
@@ 3.1.1 Resting state
In the resting state, the goal is to provide stability by using slow and less intense currents. As [1] discusses, there are ’holes’ (non-controlled ion channels) in the membrane where the counterforce is missing. The difference in the electric and thermodynamic forces could accelerate ions until the Stokes-Einstein speed (see Eq.(27)) is reached. Given that the deviation from the set-point is slight, the intensity and speed of the current are low. The slow ion transport delivers both charge and mass, so the gradients permanently direct the process control variable toward the reference point. Given that the ’resting channels’ are scattered on the surface of the membrane and it takes time for the gradients delivered by slow ions to reach the entrance of an ion channel, some slight variations in the actual value of the membrane’s potential necessarily exist. Given that the AIS comprises non-gated ion channels with much higher density than those in the membrane, the variations and the control ("leaking") are marginal, and mainly through the ion channels in the membrane’s walls, as observed in experimental physiology.
The electric voltage contribution is unidirectional; the two-component thermodynamic contributions, depending on the concentrations, can be positive or negative. If choosing the zero potential at the middle of the electric potential, we can interpret that the concentration combinations maintain their ionic concentrations (and so also thermodynamic voltage contributions) independently on both sides, starting from a zero potential level. They generate potentials autonomously in the segments proximal to the opposite surfaces of the membrane. The electric charge on the membrane creates another voltage independently. However, that voltage must fit between the two thermodynamic voltages. Any deviation in the potentials starts a current, directing the concentrations and voltages toward a balanced state. The process is complex: the sum of surface concentration of ions must be the same on both sides for the condenser, and the ratios on the intracellular and extracellular sides must be adjusted to satisfy the conditions
@@
In a balanced state, the left sides are zero: at the contact point between the electrolyte and the membrane, there are no potential differences, while in a perturbed state, they are non-zero and so they represent the driving forces for restoring the balance. (Recall that the ions are slow and that the local electric and concentration gradients control the local potentials.) From the derivation, it follows that
$U_{electric}$ $=-2*(U_{internal}^{K^{+}}$ $+U_{internal}^{Na^{+}})$ (35)
@@
Our statement is the general formulation of that "Thus ions are subject to two forces driving them across the membrane: (1) a chemical driving force, a function of the concentration gradient across the membrane, and (2) an electrical driving force, a function of the electrical potential difference across the membrane." [1], page 129 and that "nerve and muscle cells have a particularly rich variety and high density of membrane ion channels, their channels do not differ fundamentally", page 107. Fig. 9 quantitatively underpins that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH[65]. In all species, the chemical and thermodynamic processes keep balance in the cells of different construction in both resting and transient states.
According to the above derivation, we consider three characteristic segments in the neuron. In the middle, we have an electrostatically charged isolator plate that has positive and negative charge carriers on its two surfaces, and no charge carriers inside, so it has a homogeneous electric field inside it. The magnitude of the field depends only on the total concentration of the corresponding ions in the segments. On its two sides, there are ionic solutions of different compositions. If the membrane is permeable for the ions, they can move from one segment to the other as the resultant field dictates, until the fields get balanced.
@@
Figure 9: The mechanism for controlling neuronal operation is the same in all cases shown in Table 1. It appears that the thicker membrane produces a higher membrane electric potential and requires a higher thermal potential to compensate for it. However, it can solve the task using a lower salt concentration, given that the electric membrane potential accelerates the ions across the ion channels in the membrane, and the operation is faster. The electric fields for the different species have similar values, as shown in Fig. 1. Furthermore, the total concentration decreases as the membrane thickness and potential increase.
@@
The membrane’s electric potential defines the resting potential, and the "electric" and "thermodynamic" potentials work together to control neuronal operation. The anatomically defined "electric" membrane potential, plus the vast amount of ions in the extracellular segment, provides a stable concentration that serves as a solid reference point. In the resting state, the membrane potential is balanced by the presence of always-open "resting ion channels". When ions flow into the membrane, the "resting ion channels" can keep the balance: the internal $K^{+}$ and $Na^{+}$ concentrations are adjusted as the $Na^{+}$ flows in.
@@
Figure 9 also shows that in its native state, the neuron is electrically balanced: in the internal and external segments, the resulting thermodynamical and electrical forces are equal. Any foreign invasion into the neuron’s life (chemical, mechanical, or electrical, including clamping) changes the left-side reference value to an externally set value. The neuron must adjust its $Na^{+}$ and $K^{+}$ concentrations to reach a new balance while the invasion persists. That is, any invasion ("testing in the physical laboratory") changes the concentrations and voltages inside the cell. In addition, the internal biological processes can trigger internal processes that contribute (from the measurement’s point of view) foreign currents or processes, see also section 3.1.2. The feedback from the clamping methods introduces a compensating current. Even a simple conductance measurement device is active: it applies a test voltage that generates a test current, which sensitively affects the ionic composition of the neuron (recall that the concentrations may differ by up to six orders os magnitude). Although modern electronic devices attempt to conceal this effect, they subtly influence biological operation.
@@
From our result, it follows that for more chemical elements, a per-element coupled set of equations exists, plus the sum concentration agreed electric fields that are valid simultaneously and define the process variable (aka membrane potential) of the neural regulatory circuit. In a more symmetrical form
@@
$0$ $=\frac{U_{electric}^{Total}(c,\Delta z,d)}{2}+\sum_{k}U_{thermal}^{C_{k}}(d)+{invasion}\quad for\ all\ k;\ inside\ and\ outside$ (38)
@@
That is, we have an unidirectional electric potential difference and oppositely directed thermal voltages. In the balanced state, the two potentials must be equal. When changing say one concentration, it changes the other concentration on the same side, and the surface charge density of the condenser. The changed electric voltage re-adjusts the thermodynamic voltage, and as a consequence, the concentrations, in the other segment. There is no simple way to express, say, $Na^{+}]_{i}$ in the function of $K^{+}]_{i}$ or vice versa, and similarly for the other ions. Everything moves. The size matters: the amount of ions in the external segment is many orders of magnitude higher, and the gradients are reciprocally proportional to the amount of ions.
As the figure suggests, an external invasion, typically, an electric voltage on the intracellular side, changes the balanced state, and due to the parameters being linked, changes all concentrations. When moving the system out of its balanced state, in any way, a driving force appears that moves the system towards finding a new balanced state. However, recall that the ions are slow, so the changes are not instant.
Table 1 shows our results for three different famous published cases (Table 2.1 in[30]). It looks like the calibration of the concentration of negative and positive ions needs scaling, and the measurement of $Ca^{2+}$ comprises issues (the $Ca^{2+}$ concentration in all cases is inaccurate and had to be corrected).
@@
$K^{+}$ 400 20 $75.5=58log\bigl(\frac{20}{400}\bigr)$ 41.9 41.4 -62.0
$Na^{+}$ 50 440 $54.8=58log\bigl(\frac{440}{50}\bigr)$
$Cl^{-}$ 40 560 $66.5=58log\bigl(\frac{560}{40}\bigr)$ 41.9 42.0 63.0
$Ca^{2+}$ $0.27^{*}$ 10 $45.5=29log\bigl(\frac{10}{0.27}\bigr)$
@@
$K^{+}$ 124 2.25 $101=58log\bigl(\frac{2.25}{124}\bigr)$ 88.9 83.6 -125.4
$Na^{+}$ 10.4 109 $59.2=58log\bigl(\frac{109}{10.4}\bigr)$
$Cl^{-}$ 1.5 77.5 $99.4=-58log\bigl(\frac{77.5}{1.5}\bigr)$ 88.9 82.7 124.1
$Ca^{2+}$ $0.021^{*}$ 2.1 $58.0=29log\bigl(\frac{2.5}{0.25}\bigr)$
@@
$K^{+}$ 140 5 $89.7=62log\bigl(\frac{5}{140}\bigr)$ 57.7 57.3 -85.8
$Na^{+}$ 15 145 $61.1=62log\bigl(\frac{145}{15}\bigr)$
$Cl^{-}$ 4 110 $89.2=-62log\bigl(\frac{110}{4}\bigr)$ 57.7 56.9 85.9
$Ca^{2+}$ $0.04^{*}$ 5 $60.8=31log\bigl(\frac{5}{0.04}\bigr)$
@@
$Na^{+}$ 12 145 $66.6=62log\bigl(\frac{145}{12}\bigr)$ 60.7 66.6